# Estimation and goodness-of-fit testing for positive random variables with explicit Laplace transform

### v1

@@ 4 Inference for the positive stable distribution
By assuming that $X$ is distributed with the positive stable law, the corresponding Laplace transform (3) is obtained from class (4) by setting $g(x)=e^{-\lambda x}$ and $\theta=0$. Thus, the model involves two parameters $(\gamma,\lambda)$, where the parameter space is $\gamma\in\,]0,1]$ and $\lambda>0$. It is apparent that $X$ does not have a finite mean for $\gamma<1$. On the basis of (3) and (7) respectively, for $a>0$ we have
$L_{X}(a)=e^{-\lambda a^{\gamma}},\qquad{\rm and}\qquad E(Y_{a})=-{{\gamma}\over{a}}L_{X}(a)\log(L_{X}(a)),$
@@
and
$(\gamma,\lambda)=(h_{1}(m_{1,a_{*}},a_{*}),h_{2}(m_{1,a_{*}},a_{*}))=(em_{1,a_{*}}a_{*},a_{*}^{-\gamma}).$
Therefore, by using the results provided in Section 3, the estimators of $(\gamma,\lambda)$ are given by
@@
The estimators $(\widehat{\gamma},\widehat{\lambda})$ converge almost surely to $(\gamma,\lambda)$ and the random vector $\sqrt{n}(\widehat{\gamma}-\gamma,\widehat{\lambda}-\lambda)$ converges in distribution to the bivariate normal law $\mathrm{N}_{2}(0,\Sigma)$ as $n\rightarrow\infty$. Moreover, $\Sigma$ is the variance-covariance matrix of the random vector
$(S_{1},S_{2})=(a_{*}Xe^{1-a_{*}X},-\lambda e^{1-a_{*}X}(a_{*}X\log(a_{*})+1)),$
In particular, a consistent estimator of $\Sigma$ is given by the sample variance-covariance matrix of the random vectors $(\Gamma_{i},\Lambda_{i})$, with $i=1,\ldots,n$, where
$(\Gamma_{i},\Lambda_{i})=(AX_{i}e^{1-AX_{i}},-\widehat{\lambda}e^{1-AX_{i}}(X_{i}A\log(A)+1)).$
@@
$(V_{1},W)J^{\top}=(e(a_{*}V_{1}+m_{1,a_{*}}W),e\lambda\log(a_{*})(a_{*}V_{1}+m_{1,a_{*}}W)-a_{*}^{-1}\gamma\lambda W),$
@@
Since $m_{1,a_{*}}=a_{*}m_{2,a_{*}}$, by using (6) and Remark 1, we have $\widehat{L}_{X}^{(2)}(A)\sim\widehat{m}_{1,A}/A$ as $n\rightarrow\infty$ and expression (10) suggests the equivalent GOF test statistic
@@
$\sigma^{2}={\rm Var}\Big{(}e^{-a_{*}X}\Big{(}{{a_{*}m_{3,a_{*}}-2m_{2,a_{*}}}\over{m_{1,a_{*}}}}+X(1-a_{*}X)\Big{)}\Big{)}.$
In particular, $\sigma^{2}$ can be estimated by
@@
$Z_{i}=e^{-AX_{i}}\Big{(}{{A\widehat{m}_{3,A}-2\widehat{m}_{2,A}}\over{\widehat{m}_{1,A}}}+X_{i}(1-AX_{i})\Big{)},\qquad{\rm and}\qquad\overline{Z}={{1}\over{n}}\sum_{i=1}^{n}Z_{i}.$
@@
$T_{n}=\sqrt{n}((A\widehat{m}_{2,A}-\widehat{m}_{1,A})-(am_{2,a}-m_{1,a})),$
on the basis of Proposition 1 and the Delta Method, $T_{n}$ converges in distribution to the normal law $\mathrm{N}(0,\sigma^{2})$, where $\sigma^{2}$ is the variance of the r.v. $(-V_{1}+aV_{2}+m_{2,a}W)$, while $V_{1}$, $V_{2}$ and $W$ are defined in Proposition 1. It holds
$-V_{1}+a_{*}V_{2}+m_{2,a_{*}}W=e^{-a_{*}X}\Big{(}-X(1-a_{*}X)-{{a_{*}m_{3,a_{*}}}\over{m_{1,a_{*}}}}+2{{m_{2,a_{*}}}\over{m_{1,a_{*}}}}\Big{)},$
@@ 5 Inference for the Tweedie distribution
By assuming that $X$ is distributed with the Tweedie law, the corresponding Laplace transform (5) is a special case of the class (4) by setting
@@
$a_{*}=\Big{(}\frac{1}{{\rm sgn}(\gamma)\lambda}+\theta^{\gamma}\Big{)}^{1/\gamma}-\theta.$
@@
$E(Y^{3}_{a_{*}})={{E(Y_{a_{*}})^{3}}\over{L^{2}_{X}(a_{*})}}+{{E(Y_{a_{*}})(1-\gamma)}\over{\theta+a_{*}}}\Big{(}{{3E(Y_{a_{*}})}\over{L_{X}(a_{*})}}+{{2-\gamma}\over{\theta+a_{*}}}\Big{)}.$
@@
$\kern-0.85358pt\gamma\kern-1.99168pt=\kern-1.99168pth_{1}(m_{1,a_{*}}\kern-0.28453pt,\kern-0.28453ptm_{2,a_{*}}\kern-0.28453pt,\kern-0.28453ptm_{3,a_{*}}\kern-0.28453pt,\kern-0.28453pta_{*})\kern-1.99168pt=\kern-1.99168pt1\kern-0.85358pt-\kern-0.85358pt\big{(}{{m_{2,a_{*}}}\over{m_{1,a_{*}}^{2}}}\kern-0.85358pt-\kern-0.85358pte\big{)}\phi(m_{1,a_{*}}\kern-0.28453pt,\kern-0.28453ptm_{2,a_{*}},m_{3,a_{*}}),$
@@
$\widehat{\gamma}=1-\Big{(}{{\widehat{m}_{2,A}}\over{\widehat{m}_{1,A}^{2}}}-e\Big{)}\phi(\widehat{m}_{1,A},\widehat{m}_{2,A},\widehat{m}_{3,A}),$
$\widehat{\lambda}=e\widehat{m}_{1,A}|\widehat{\gamma}|^{-1}(\widehat{\theta}+A)^{1-\widehat{\gamma}},\qquad\widehat{\theta}=-A+\widehat{m}_{1,A}^{-1}\,\phi(\widehat{m}_{1,A},\widehat{m}_{2,A},\widehat{m}_{3,A}).$
@@
If $\lambda\theta^{\gamma}>-{\rm sgn}(\gamma)$, the estimators $(\widehat{\gamma},\widehat{\lambda},\widehat{\theta})$ converges almost surely to $(\gamma,\lambda,\theta)$ and $\sqrt{n}(\widehat{\gamma}-\gamma,\widehat{\lambda}-\lambda,\widehat{\theta}-\theta)$ converges in distribution to the trivariate normal law $\mathrm{N}_{3}(0,\Sigma)$ as $n\rightarrow\infty$. Moreover, $\Sigma$ is the variance-covariance matrix of the random vector
@@
$T^{\prime}_{n}=\sqrt{n}({\rm sgn}(\widehat{\gamma})\widehat{\lambda}((\widehat{\theta}+A)^{\widehat{\gamma}}-\widehat{\theta}^{\,\widehat{\gamma}})-1).$
The previous test statistic has the same asymptotic behaviour of
$\widetilde{T}^{\prime}_{n}=\sqrt{n}\Big{(}1-\Big{(}{{\widehat{\theta}}\over{\widehat{\theta}+A}}\Big{)}^{\widehat{\gamma}}-{{\widehat{\gamma}}\over{e\widehat{m}_{1,A}(\widehat{\theta}+A)}}\Big{)}.$ (14)
The large-sample properties of (14) are derived in the following Proposition.
@@
$\psi=\psi(m_{1},m_{2},m_{3})={{m_{3}-e^{2}m_{1}^{3}}\over{m_{1}m_{2}-em_{1}^{3}}}-2e-{{m_{2}}\over{m_{1}^{2}}}$
and
$\phi=\phi(m_{1},m_{2},m_{3})=1-{{1}\over{\psi}}\big{(}{{m_{2}}\over{m_{1}^{2}}}-e\big{)}.$
@@
$\widetilde{T}^{\prime}_{n}=\sqrt{n}\Big{(}-(1-A\widehat{m}_{1,A}\widehat{\psi})^{\widehat{\phi}}-e^{-1}\Big{(}\widehat{\psi}-{{\widehat{m}_{2,A}}\over{\widehat{m}_{1,A}^{2}}}\Big{)}\Big{)}$
@@
$\widehat{\psi}={{\widehat{m}_{3,A}-e^{2}\widehat{m}_{1,A}^{3}}\over{\widehat{m}_{1,A}\widehat{m}_{2,A}-e\widehat{m}_{1,A}^{3}}}-2e-{{\widehat{m}_{2,A}}\over{\widehat{m}_{1,A}^{2}}}$
and $\widehat{\phi}=1-{{1}\over{\widehat{\psi}}}\big{(}{{\widehat{m}_{2,A}}\over{\widehat{m}_{1,A}^{2}}}-e\big{)},$ on the basis of Proposition 1 and of the Delta Method, by considering the equality
@@ 6 Simulation results
We assess the performance of the estimators and GOF procedures proposed in Section 4 and Section 5 for the positive stable and Tweedie distributions, respectively. The aim is to inspect their finite sample behaviour by means of a Monte Carlo simulation study. In particular, all the numerical experiments that we consider are carried out on the basis of 3500 independent replicas.
In the following we denote the positive stable law as $\mathrm{PS}(\gamma,\lambda)$. Moreover, since we also consider the more complex situation for which $\gamma<0$, i.e. when the Tweedie distribution is a compound Poisson law, we adopt the notation with a null subscript $\mathrm{TW}_{0}(\mu,w,p)$ for the Tweedie law with the parameterization
$\mu=|\gamma|\lambda\theta^{\gamma-1},\qquad w=\frac{1-\gamma}{\theta},\qquad p=\exp(-\lambda\theta^{\gamma}).$ (15)
In this case, $\mu$ and $p$ respectively represent the mean and the probability of observing a zero value. The standard parameterization is denoted by $\mathrm{TW}(\gamma,\lambda,\theta)$.
We begin with the assessment of the proposed estimators performance in terms of Relative Root Mean Square Error (RRMSE), obtained as the root mean square error divided by the true parameter value. The results, reported in Table 1 for the positive stable law and in Table 2 for the Tweedie law, show consistency of the estimators for all the considered parameter values. Moreover, Figures 1 and 2 highlight that the empirical coverage of the interval estimators for $(\gamma,\lambda)$ of the positive stable distribution attains the nominal level of $0.95$ for moderately large sample sizes $n=100,200$.
@@
Figure 1: Empirical coverage for $\gamma$ of the positive stable distribution (nominal level 0.95). Dashed blue line ($n=100$). Solid red line ($n=200$).
Figure 2: Empirical coverage for $\lambda$ of the positive stable distribution (nominal level 0.95). Dashed blue line ($n=100$). Solid red line ($n=200$).
Concerning the GOF, both the testing procedures based on (13) and (14) are evaluated in terms of the actual significance and power levels. To evaluate the actual significance of the two tests, independent Monte Carlo samples were generated under the null hypothesis $H_{0}$, which represents the functional hypothesis that the target r.v. is distributed according to the positive stable distribution or the Tweedie distribution. The random variate generation was carried out by considering the stochastic representations discussed in Devroye and James, (2014) and Barabesi et al., 2016b .
First, we obtained the empirical actual significance level of (13) under the positive stable model for various representative parameter choices. The results are reported as percentage of rejections at nominal level equal to $0.05$ for $n=100,200,300$ (see Table 3). The significance level was also depicted in Figure 3 for varying $\lambda$ by assuming $\gamma=0.3,0.5,0.7,0.8$ and for $n=100,200$. From Table 3 and Figure 3, it is apparent that the test shows a satisfactory actual significance level, approaching the nominal level as $n$ increases, albeit it is slightly conservative for small $\gamma$ when the sample size is small.
Table 3: Percentage values of the empirical significance level (nominal level fixed to 0.05) of the positive stable GOF test (13) for some parameter choices.
@@
Figure 3: Empirical significance level (nominal level fixed to 0.05) with $n=100$ (dashed line) and $n=200$ (solid line).
Subsequently, we computed the empirical actual significance level of (14) under the Tweedie model for some parameter choices under both the alternative parameterization (15) and the standard one (see Table 7 for the correspondence between the parameter values). The results are reported as percentage of rejections at nominal level $0.05$ for $n=300,500,1000,1500$ (see Table 4). As it could be expected, the testing problem is much more challenging for the Tweedie GOF test owing to the additional parameter. Indeed, the test achieves the nominal level for a larger sample size such as $n=1500$. Anyway, it should be remarked that the test is conservative.
Table 4: Percentage values of the empirical significance level (nominal level fixed to 0.05) of the Tweedie GOF test (14) for some parameter choices.
@@
As to the power of the two tests, we considered as alternative distributions the Linnik law denoted by $\mathrm{LI}(\gamma,\lambda,\delta)$, the shape-scale Pareto law denoted by $\mathrm{PA}(\alpha,\beta)$ with density function $f(x)=\alpha\beta^{\alpha}/x^{\alpha+1}\mathbf{1}_{\{x\geq\beta\}}$, and the shape-scale Weibull law denoted by $\mathrm{WE}(k,\lambda)$ with density function $f(x)=(k/\lambda)(x/\lambda)^{k-1}\exp(-(x/\lambda)^{k})$. Moreover, we considered the log-normal law denoted by $\mathrm{LN}(\mu,\sigma)$ and, as a further alternative, $\mathrm{LN}^{1/2}(\mu,\sigma)$, that represent respectively the laws of the r.v.s $Y=e^{X}$ and $Y=e^{X^{2}}$, where $X$ has normal law $\mathrm{N}(\mu,\sigma)$. We also considered some zero-inflated distributions, where the third parameter represents the probability of observing a zero value. Such laws are indicated by adopting a null subscript, e.g. $\mathrm{PA}_{0}(\alpha,\beta,p)$. From Table 5, the positive stable GOF show an appreciable power for all the considered alternatives. In particular, it achieves an elevate power for the $\mathrm{PA}$ and $\mathrm{LN}^{1/2}$ distribution. In contrast $\mathrm{LI}$ is similar to $\mathrm{PS}$, justifying a smaller power. The results in Table 6 evidence that the Tweedie GOF test is consistent, even if the increase of the number of parameters leads to a more challenging testing problem in terms of type 2 error. Indeed, larger samples are necessary to achieve a power comparable to the positive stable case, even if the power is satisfactory for $n$ larger than $500$. Moreover, under zero-inflated models, the procedure achieves an increasing performance as the zero probability increases.
Table 5: Percentage values of the empirical power of the positive stable GOF test (13) for some selected alternative distributions.
@@
Table 6: Percentage values of the empirical power of the Tweedie GOF test (14) for some selected alternative distributions.
@@
Table 7: Corresponding values of $(\gamma,\lambda,\theta)$ for the alternative parameterization $\mathrm{TW}_{0}(\mu,w,p)$ given in (15) for the Tweedie distribution.

### v2

@@ 4 Inference for the positive stable distribution
In this section we assume that $X$ is distributed with a positive stable law. The corresponding Laplace transform (3) is obtained from class (4) by setting $g(x)=e^{-\lambda x}$ and $\theta=0$. Thus, the model involves two parameters $(\gamma,\lambda)$, where the parameter space is $\gamma\in\,]0,1]$ and $\lambda>0$. It is apparent that $X$ does not have a finite mean for $\gamma<1$. On the basis of (3) and (7) respectively, for $a>0$ we have
$L_{X}(a)=e^{-\lambda a^{\gamma}},\qquad{\rm and}\qquad E(Y_{a})=E(Xe^{-aX})=-{{\gamma}\over{a}}L_{X}(a)\log(L_{X}(a)),$
@@
and $(\gamma,\lambda)=(h_{1}(m_{1,a_{*}},a_{*}),h_{2}(m_{1,a_{*}},a_{*}))=(em_{1,a_{*}}a_{*},a_{*}^{-\gamma}).$ Therefore, by using the results provided in Section 3, the estimators of $(\gamma,\lambda)$ are given by
@@
The estimators $(\widehat{\gamma},\widehat{\lambda})$ converge almost surely to $(\gamma,\lambda)$ and the random vector $\sqrt{n}(\widehat{\gamma}-\gamma,\widehat{\lambda}-\lambda)$ converges in distribution to the bivariate normal law $\mathrm{N}_{2}(0,\Sigma)$ as $n\rightarrow\infty$. Here, $\Sigma$ is the variance-covariance matrix of the random vector
$(S_{1},S_{2})=(a_{*}Xe^{1-a_{*}X},-\lambda e^{1-a_{*}X}\{a_{*}X\log(a_{*})+1\})$
A consistent estimator of $\Sigma$ is given by the sample variance-covariance matrix of the random vectors $(\Gamma_{i},\Lambda_{i})$, with $i=1,\ldots,n$, where
$(\Gamma_{i},\Lambda_{i})=(AX_{i}e^{1-AX_{i}},-\widehat{\lambda}e^{1-AX_{i}}\{X_{i}A\log(A)+1\}).$
@@
$(V_{1},W)J^{\top}=(e\{a_{*}V_{1}+m_{1,a_{*}}W\},e\lambda\log(a_{*})\{a_{*}V_{1}+m_{1,a_{*}}W\}-a_{*}^{-1}\gamma\lambda W),$
@@
Since $m_{1,a_{*}}=a_{*}m_{2,a_{*}}$, by using (6), we have $\widehat{L}_{X}^{(2)}(A)\sim\widehat{m}_{1,A}/A$ as $n\rightarrow\infty$ and expression (8) suggests the equivalent GOF test statistic
@@
$\sigma^{2}={\rm Var}\Big{(}e^{-a_{*}X}\Big{\{}{{a_{*}m_{3,a_{*}}-2m_{2,a_{*}}}\over{m_{1,a_{*}}}}+X(1-a_{*}X)\Big{\}}\Big{)}.$
$\sigma^{2}$ can be estimated by
@@
$Z_{i}=e^{-AX_{i}}\Big{\{}{{A\widehat{m}_{3,A}-2\widehat{m}_{2,A}}\over{\widehat{m}_{1,A}}}+X_{i}(1-AX_{i})\Big{\}},\qquad{\rm and}\qquad\overline{Z}={{1}\over{n}}\sum_{i=1}^{n}Z_{i}.$
@@
$T_{n}=\sqrt{n}\{(A\widehat{m}_{2,A}-\widehat{m}_{1,A})-(am_{2,a}-m_{1,a})\},$
on the basis of Proposition 1 and the Delta Method, $T_{n}$ converges in distribution to the normal law $\mathrm{N}(0,\sigma^{2})$, where $\sigma^{2}$ is the variance of the r.v. $(-V_{1}+aV_{2}+m_{2,a}W)$, where $V_{1}$, $V_{2}$ and $W$ are defined in Proposition 1. It holds
$-V_{1}+a_{*}V_{2}+m_{2,a_{*}}W=e^{-a_{*}X}\Big{\{}-X(1-a_{*}X)-{{a_{*}m_{3,a_{*}}}\over{m_{1,a_{*}}}}+2{{m_{2,a_{*}}}\over{m_{1,a_{*}}}}\Big{\}},$
@@ 5 Inference for the Tweedie distribution
In this section we assume that $X$ is distributed with the Tweedie law. The corresponding Laplace transform (5) is a special case of the class (4) by setting
@@
$a_{*}=\Big{\{}\frac{1}{{\rm sgn}(\gamma)\lambda}+\theta^{\gamma}\Big{\}}^{1/\gamma}-\theta.$
@@
$E(Y^{3}_{a_{*}})={{E(Y_{a_{*}})^{3}}\over{L^{2}_{X}(a_{*})}}+{{E(Y_{a_{*}})(1-\gamma)}\over{\theta+a_{*}}}\Big{\{}{{3E(Y_{a_{*}})}\over{L_{X}(a_{*})}}+{{2-\gamma}\over{\theta+a_{*}}}\Big{\}}.$
@@
$\gamma=h_{1}(m_{1,a_{*}},m_{2,a_{*}},m_{3,a_{*}},a_{*})=1-\Big{(}{{m_{2,a_{*}}}\over{m_{1,a_{*}}^{2}}}-e\Big{)}\phi(m_{1,a_{*}},m_{2,a_{*}},m_{3,a_{*}}),$
@@
$\widehat{\gamma}$ $=1-\Big{(}{{\widehat{m}_{2,A}}\over{\widehat{m}_{1,A}^{2}}}-e\Big{)}\phi(\widehat{m}_{1,A},\widehat{m}_{2,A},\widehat{m}_{3,A}),$ (14)
$\widehat{\lambda}$ $=e\widehat{m}_{1,A}|\widehat{\gamma}|^{-1}(\widehat{\theta}+A)^{1-\widehat{\gamma}},$ (15)
$\widehat{\theta}$ $=-A+\widehat{m}_{1,A}^{-1}\,\phi(\widehat{m}_{1,A},\widehat{m}_{2,A},\widehat{m}_{3,A}).$ (16)
@@
If $\lambda\theta^{\gamma}>-{\rm sgn}(\gamma)$, the estimators $(\widehat{\gamma},\widehat{\lambda},\widehat{\theta})$ converges almost surely to $(\gamma,\lambda,\theta)$ and $\sqrt{n}(\widehat{\gamma}-\gamma,\widehat{\lambda}-\lambda,\widehat{\theta}-\theta)$ converges in distribution to the trivariate normal law $\mathrm{N}_{3}(0,\Sigma)$ as $n\rightarrow\infty$. Here, $\Sigma$ is the variance-covariance matrix of the random vector
@@
$T^{\prime}_{n}=\sqrt{n}[{\rm sgn}(\widehat{\gamma})\widehat{\lambda}\{(\widehat{\theta}+A)^{\widehat{\gamma}}-\widehat{\theta}^{\,\widehat{\gamma}}\}-1].$
The previous test statistic, suitably normalized, has the same asymptotic behaviour of the normalized version of
$\widetilde{T}^{\prime}_{n}=\sqrt{n}\Big{\{}1-\Big{(}{{\widehat{\theta}}\over{\widehat{\theta}+A}}\Big{)}^{\widehat{\gamma}}-{{\widehat{\gamma}}\over{e\widehat{m}_{1,A}(\widehat{\theta}+A)}}\Big{\}}.$ (17)
The large-sample properties of the test statistic $\widetilde{T}^{\prime}_{n}$ in (17) are derived in the following Proposition.
@@
$\psi=\psi(m_{1},m_{2},m_{3})={{m_{3}-e^{2}m_{1}^{3}}\over{m_{1}m_{2}-em_{1}^{3}}}-2e-{{m_{2}}\over{m_{1}^{2}}},\quad\phi=\phi(m_{1},m_{2},m_{3})=1-{{1}\over{\psi}}\Big{(}{{m_{2}}\over{m_{1}^{2}}}-e\Big{)}.$
@@
$\widetilde{T}^{\prime}_{n}=\sqrt{n}\Big{\{}-(1-A\widehat{m}_{1,A}\widehat{\psi})^{\widehat{\phi}}-e^{-1}\Big{(}\widehat{\psi}-{{\widehat{m}_{2,A}}\over{\widehat{m}_{1,A}^{2}}}\Big{)}\Big{\}}$
@@
$\widehat{\psi}={{\widehat{m}_{3,A}-e^{2}\widehat{m}_{1,A}^{3}}\over{\widehat{m}_{1,A}\widehat{m}_{2,A}-e\widehat{m}_{1,A}^{3}}}-2e-{{\widehat{m}_{2,A}}\over{\widehat{m}_{1,A}^{2}}},\ \quad\ \widehat{\phi}=1-{{1}\over{\widehat{\psi}}}\Big{(}{{\widehat{m}_{2,A}}\over{\widehat{m}_{1,A}^{2}}}-e\Big{)},$
on the basis of Proposition 1 and of the Delta Method, by considering the equality
@@ 6 Simulation results
By means of a Monte Carlo simulation study, we assess the finite sample performance of the estimators and GOF procedures proposed in Section 4 for the positive stable PS$(\gamma,\lambda)$ distribution and in Section 5 for the Tweedie TW$(\gamma,\lambda,\theta)$ distribution. For the Tweedie distribution, when it is a compound Poisson law, that is when $\gamma<0$, it can be useful to determine the parameters values characterizing the distributions from which data are generated using the following notation
${\rm TW}(\gamma,\lambda,\theta)={\rm TW}_{0}(|\gamma|\lambda\theta^{\gamma-1},\frac{1-\gamma}{\theta},\exp(-\lambda\theta^{\gamma}))={\rm TW}_{0}(\mu,w,p).$ (18)
It should be noted that $\mu$ and $p$ have an immediate interpretation as they represent the mean and the probability of observing a zero value, respectively. Then, if $\gamma<0$, the model parameters are $(\gamma,\lambda,\theta)$, but their values are derived from the values selected for $\mu$, $w$ and $p$. The random variate generation is carried out by considering the stochastic representations discussed in Devroye and James, (2014) and Barabesi et al., 2016b .
As to the positive stable distribution, first we consider four choices of the parameter values and, for each choice, $3500$ samples of size $n=100,200,300$ are independently generated. For each sample, parameter estimates are obtained by means of (12). From the Monte Carlo distributions of the estimates, the Relative Root Mean Squared Error (RRMSE) of $\widehat{\gamma}$ and $\widehat{\lambda}$ is computed as the root mean square error divided by the true parameter value. Analogously, two parameter choices are considered for each parametrization of the Tweedie distribution and, for each choice, $3500$ samples of size $n=500,1000,1500$ are independently generated. The different sample sizes considered for the two distributions are related to the different number of parameters. For each sample the estimates of $\gamma$, $\lambda$ and $\theta$ are obtained by means of (14), (15) and (16), and the RRMSE values are computed from the corresponding Monte Carlo distributions. Percentage values of RRMSEs are reported in Table 1 and in Table 2 for the positive stable law and for the Tweedie law, respectively. In these Tables and in all subsequent ones, subscripts denote the sample size. The results illustrate the consistency of the estimators for all the considered parameter values.
Moreover, for the positive stable distribution, also confidence interval estimates at confidence level $0.951-\alpha=$ 0.95 are obtained using the quantiles of the standard normal distribution and results in Section 4. The empirical coverage is computed as the proportion of times the true value is in the interval. More precisely, for any fixed value $\gamma=0.3,0.5,0.7,0.8$ and for $\lambda$ varying from $0.5$ to $12$ by $0.5$, $3500$ independent samples of $n=100,200$ are generated and the empirical coverages of the 0.95 confidence intervals are computed for $\gamma$ and $\lambda$ and depicted in Figure 1 and Figure 2, respectively. The figures highlight that the empirical coverage of the interval estimators for both parameters attains the nominal level of $0.95$ for moderately large sample sizes.
@@
Figure 1: Empirical coverage of the 0.95 confidence interval for $\gamma$ in positive stable distributions for different choices of $\gamma$ and $\lambda$. Dashed line for $n=100$ and solid line for $n=200$.
Figure 2: Empirical coverage of the 0.95 confidence interval for $\lambda$ in positive stable distributions for different choices of $\gamma$ and $\lambda$. Dashed line for $n=100$ and solid line for $n=200$.
Concerning the GOF, both the testing procedures based on (13) and (17) are evaluated in terms of the empirical significance and power levels. First, $3500$ independent samples are generated under the null hypothesis $H_{0}$, which represents the functional hypothesis that the target r.v. is distributed according to the positive stable distribution or the Tweedie distribution. Then, the empirical significance level is computed as the proportion of rejections at the nominal level equal to $0.05$.
Under the positive stable model, we obtain the empirical significance level for various representative parameter choices and for samples of size $n=100,200,300$ (see Table 3). The empirical significance level is also depicted in Figure 3 for $\gamma=0.3,0.5,0.7,0.8$ and $\lambda$ varying from $0.5$ to $12$ by $0.5$, for $n=100,200$. From Table 3 and Figure 3, it is apparent that the test shows a satisfactory significance level, approaching the nominal level as $n$ increases, albeit it is slightly conservative for small $\gamma$ when the sample size is small. Analogously, under the Tweedie model, we compute the empirical significance level for some parameter choices for $n=500,1000,1500$. The parameter values when $\gamma<0$ are given in Table 7. The results are reported in Table 4. As it could be expected, the testing problem is much more challenging for the Tweedie GOF test owing to the additional parameter. Indeed, the test achieves the nominal level for a larger sample size such as $n=1500$. Anyway, it should be remarked that the test is conservative.
Table 3: Percentage values of the empirical significance level (nominal level 0.05) of the positive stable GOF test for some parameter choices.
@@
Figure 3: Empirical significance level (nominal level 0.05) for different choices of $\gamma$ and $\lambda$. Dashed line for $n=100$ and solid line for $n=200$.
Table 4: Percentage values of the empirical significance level (nominal level 0.05) of the Tweedie GOF test for some parameter choices.
@@
As to the power of the two tests, we consider as alternative distributions the Linnik law denoted by $\mathrm{LI}(\gamma,\lambda,\delta)$, the shape-scale Pareto law denoted by $\mathrm{PA}(\alpha,\beta)$ with density function $f(x)=\alpha\beta^{\alpha}/x^{\alpha+1}\mathbf{1}_{\{x\geq\beta\}}$, and the shape-scale Weibull law denoted by $\mathrm{WE}(k,\lambda)$ with density function $f(x)=(k/\lambda)(x/\lambda)^{k-1}\exp(-(x/\lambda)^{k})$. Moreover, we consider the log-normal law denoted by $\mathrm{LN}(\mu,\sigma)$ and, as a further alternative, $\mathrm{LN}^{1/2}(\mu,\sigma)$, that represent respectively the laws of the r.v.s $Y=e^{X}$ and $Y=e^{X^{2}}$, where $X$ has normal law $\mathrm{N}(\mu,\sigma)$. We also consider some zero-inflated distributions, where the third parameter represents the probability of observing a zero value. Such laws are indicated by adopting a null subscript, e.g. $\mathrm{PA}_{0}(\alpha,\beta,p)$. The empirical powers of the tests, computed as the proportion of rejections when data are generated according to alternative distributions, are reported in Table 5 and Table 6 under the null hypothesis of positive stable or Tweedie distributions, respectively. In particular, for each alternative distribution, we generate $3500$ independent samples of size $n=100,200,300$ when performing GOF test at significance level $0.05$ for the positive stable distribution and of size $n=300,500,1000,1500$ for the Tweedie distribution.
From Table 5, the positive stable GOF show an appreciable power for all the considered alternatives. In particular, it achieves an elevate power for the $\mathrm{PA}$ and $\mathrm{LN}^{1/2}$ distribution. In contrast $\mathrm{LI}$ is similar to $\mathrm{PS}$, justifying a smaller power. The results in Table 6 evidence that the Tweedie GOF test is consistent, even if the increase in the number of parameters leads to a more challenging testing problem in terms of type 2 error. Indeed, larger samples are necessary to achieve a power comparable to the positive stable case, even if the power is satisfactory for $n$ larger than $500$. Moreover, under zero-inflated models, the procedure achieves an increasing performance as the zero probability increases.
Table 5: Percentage values of the empirical power of the positive stable GOF test for some selected alternative distributions.
@@
Table 6: Percentage values of the empirical power of the Tweedie GOF test for some selected alternative distributions.
@@
Table 7: Correspondence between the values of $(\gamma,\lambda,\theta)$ and $(\mu,w,p)$ for the Tweedie distribution from (18).